2x – y = 6
-x + y = 2

Answer by stanbon(75887)

2x – y = 6
-x + y = 2
-----
Add and solve for "x":
x = 8
----
Solve for "y":
2x - y = 6
16 - y = 6
y = 10
------------
Cheers,
Stan H.